# STP: Self-play LLM Theorem Provers with Iterative Conjecturing and Proving

### v2

@@ A.9 Compute resources
Our experiments are primarily done on TPU-v4 VMs with 32 nodes. Each node contains 4 TPU chips (8 TPU cores), 240 CPU cores, and 400G memory. We use vLLM [Kwon et al., 2023] to generate LLM outputs, and Levanter [^8] https://github.com/stanford-crfm/levanter to train the LLM. In both STP and expert iteration, since the generated proofs are heavily filtered (based on the correctness, elegancy, trivialness, etc.) when constructing the training dataset, the LLM training only takes less than 25% of the wall-clock time for TPU compute, and generating proofs takes the rest 75%. In the Lean experiments, running the verifier requires 2x wall-clock time than the TPU compute. Therefore, even though we parallelize the CPU and TPU compute, the TPU is still idle for more than 50% of the wall-clock time.
@@ B.1 Additional results with Lean
In Table 3, we compare the performance of our method with prior works on PutnamBench. Note that DSP [Jiang et al., 2022a] uses Isabelle verifier where PutnamBench only has 640 statements. Our model STP achieves state-of-the-art performance by solving 9 out of 644 problems.
Table 2 compares the model obtained by final re-training with and without the proofs of generated conjectures, as discussed in the ablation study section (Section 4.4). The results show that it is still beneficial to re-train with the generated conjectures in addition to the successfully proved statements in LeanWorkBook even for performance on miniF2F-test and ProofNet-tess, which it leads to about 1% performance gain (for pass@128).
@@ B.2 Additional results with Isabelle
In Fig. 5 (Right), we plot the pass rates of STP and baseline methods on LeanWorkbook starting from iteration 0. The red crosses shows the points where we refresh the training process as described in Section 4.1.
Table 3: Results on PutnamBench [Tsoukalas et al., 2024].

### v4

@@ A.9 Compute resources
Our experiments are primarily done on TPU-v4 VMs with 32 nodes. Each node contains 4 TPU chips (8 TPU cores), 240 CPU cores, and 400G memory. We use vLLM [Kwon et al., 2023] to generate LLM outputs, and Levanter [^9] https://github.com/stanford-crfm/levanter to train the LLM. In both STP and expert iteration, since the generated proofs are heavily filtered (based on the correctness, elegancy, trivialness, etc.) when constructing the training dataset, LLM training only takes less than 25% of the wall-clock time for TPU compute, and generating proofs takes the rest 75%.
@@ B.1 Additional results with Lean
In Table 3, we compare the performance of our method with prior works on PutnamBench. Note that DSP [Jiang et al., 2022a] uses Isabelle verifier where PutnamBench only has 640 statements. Our model STP achieves state-of-the-art performance by solving 8 out of 644 problems.
Table 2 compares the model obtained by final re-training with and without the proofs of generated conjectures, as discussed in the ablation study section (Section 4.4). The results show that it is still beneficial to re-train with the generated conjectures in addition to the successfully proved statements in LeanWorkBook even for performance on miniF2F-test and ProofNet-test, which leads to about 2-3% performance gain (for pass@128).
@@ B.2 Additional results with Isabelle
In Fig. 5 (Right), we plot the pass rates of STP and baseline methods on LeanWorkbook starting from iteration 0. The red crosses shows the points where we refresh the training process as described in Section 4.1. Our models are tested with PutnamBench [Tsoukalas et al., 2024], commit d49896f. [^10] https://github.com/trishullab/PutnamBench/tree/d49896fdc87a128a70e15a185d8dfca3516dd894
Table 3: Results on PutnamBench.
@@
B.3 Examples of unproved statements in LeanWorkbook
In this section, we list 20 randomly selected statements from LeanWorkbook that are not proved during STP training. The following table shows the formal statement, the corresponding natural language statement in LeanWorkbook, the correctness of formalization, and the correctness of the formal statement.
{xltabular}
|c|X|X|p1.7cm|p1.5cm| # Lean formal statement Natural language statement Correct Correct
formalization? statement?
1 theorem lean_workbook_7116 (x y z : ℝ) (hx : x + y + z = 3) : x ^ 2 + y ^ 2 + %z ^ 2 + 3 2 * (1 / x ^ 2 + 1 / y ^ 2 + 1 / z ^ 2) := by If $a=x^{2},b=y^{2},c=z^{2}$ it suffices to show that: $x+y+z=3\implies x^{2}+y^{2}+z^{2}+3\leq 2(\frac{1}{x^{2}}+\frac{1}{y^{2}}+\frac{1}{z^{2}})$ Yes No. The case $x=0$ is ill defined.
2 theorem lean_workbook_plus_72390 (a b n : ℕ) (h : a b [ZMOD n]) : a^n b^n %[ZMOD n^2] := by Prove that if $a\equiv b\pmod{n}$, then $a^{n}\equiv b^{n}\pmod{n^{2}}$. Yes Yes
3 theorem lean_workbook_35349 (a b c : ℝ) : (9 / (a + b + c + Real.sqrt (3 * (a %* b + b * c + c * a)))) (1 / (a + b) + 1 / (b + c) + 1 / (c + a)) := by For: $\frac{9}{a+b+c+\sqrt{3(ab+bc+ca)}}\leq\frac{1}{a+b}+\frac{1}{b+c}+\frac{1}{c+a}$ Yes (but maybe missing the implicit assumption $a,b,c>0$) No (e.g., $(a,b,c)=(-0.5,1,1)$.)
4 theorem lean_workbook_8880 (a b c : ℝ) : a * Real.sqrt (b ^ 2 + c ^ 2) + b * R%eal.sqrt (c ^ 2 + a ^ 2) + c * Real.sqrt (a ^ 2 + b ^ 2) 3 * Real.sqrt 2 :=by Prove that $a\sqrt{b^{2}+c^{2}}+b\sqrt{c^{2}+a^{2}}+c\sqrt{a^{2}+b^{2}}\leq 3\sqrt{2},$ Yes No
5 theorem lean_workbook_plus_44018 (x : ℝ) (hx : 0 < x) (a : ℝ) (ha : a = x^(1/3%)) : a^2 - 2*a - (a^3 - 4)*Real.sqrt (a^3 - 7) - 3*a^3 + 28 = 0 := by Put $3]{x}=a$ The equation is equivalent to $a^{2}-2a-(a^{3}-4)\sqrt{a^{3}-7}-3a^{3}+28=0$ Yes No
6 theorem lean_workbook_plus_35882 (a b c : ℝ) (ha : 0 < a) (hb : 0 < b) (hc : 0% < c) : (Real.sqrt ((a + 2 * b + 3 * c) / (4 * a + b + c)) + Real.sqrt ((3 * a+ b + 2 * c) / (a + 4 * b + c)) + Real.sqrt ((2 * a + 3 * b + c) / (a + b + 4% * c))) 3 := by If $a,b,c>0$ prove or disprove $\sqrt{\frac{a+2b+3c}{4a+b+c}}+\sqrt{\frac{3a+b+2c}{a+4b+c}}+\sqrt{\frac{2a+3b+c}{a+b+4c}}\geq 3$ Yes Yes
7 theorem lean_workbook_12619 : x y : ℝ, (x^2+x+xy+y^2) 1 -(1/3)*Real.sqr%t ((1/2)*(69+11*Real.sqrt 33)) x^2+2*x*y x^2+2*x*y (1/3)*Real.sqrt ((1/2%)*(69+11*Real.sqrt 33)) := by Let $x^{2}+x+xy+y^{2}\leq 1(x,y\in R)$ Prove that $-\frac{1}{3}\sqrt{\frac{1}{2}(69+11\sqrt{33})}\leq x^{2}+2xy\leq\frac{1}{3}\sqrt{\frac{1}{2}(69+11\sqrt{33})}$ No (there is a xy term in Lean. Should be x * y) No
8 theorem lean_workbook_plus_20629 (f : ℝ) (x : ℝ) : f (f x + 1) = f x + 1 % := by Prove that $f(f(x)+1)=f(x)+1$ for all real $x$. Yes No
9 theorem lean_workbook_37208 (n : ℕ) (hn : 0 < n) : (n : ℝ) / (n! : ℝ) ^ (1 / n%) < (1 + 1 / n)^n := by Prove that: ${\frac{n}{\sqrt[n]{n!}}<\Big{(}1+\frac{1}{n}\Big{)}^{n}}$ for every positive integer $n$ Yes Yes
10 theorem lean_workbook_10259 (a b : ℕ) (hab : a b) (h : a + b $\mid$ a^2 + b^2) : a * b + 4 (Nat.gcd a b)^4 := by Given $a,b\in\mathbb{N}(a\neq b)$ so that $a+b\mid a^{2}+b^{2}$ Let $d=gcd(a,b)$ Prove that $ab+4\leq d^{4}$ Yes Yes
11 theorem lean_workbook_45322 (a b : ℝ) (ha : 0 < a) (hb : 0 < b) (hab : (a + 1 %/ a) * (b + 1 / b) = 2 + 3 / Real.sqrt 2) : 1 a ^ 4 + b ^ 4 a ^ 4 + b ^ 4 %≤ 4 := by Let $a,b>0$ and $(a+\frac{1}{a})(b+\frac{1}{b})=2+\frac{3}{\sqrt{2}}.$ Prove that $1\leq a^{4}+b^{4}\leq 4$ Yes Yes
12 theorem lean_workbook_28189 (x y z : ℝ) : Real.sqrt (1 + 48 * x / (y + z)) (%184 * x ^ 2 - 32 * (y ^ 2 + z ^ 2) + 289 * x * (y + z) + 127 * y * z) / (8 * (%x ^ 2 + y ^ 2 + z ^ 2) + 47 * (y * z + z * x + x * y)) := by prove: $\sqrt{1+\frac{48x}{y+z}}\geq\frac{184x^{2}-32(y^{2}+z^{2})+289x(y+z)+127yz}{8(x^{2}+y^{2}+z^{2})+47(yz+zx+xy)}$ Yes (but maybe missing $x\geq 0$) No $x=-1/24,y=z=1$
13 theorem lean_workbook_31673 (x : ℝ) (h : ∑’ k : ℕ, (7 / (2^k)) = x) : x %= 14 := by Solution without Geometric Formula71+72+74+$\frac{7}{1}+\frac{7}{2}+\frac{7}{4}+\frac{7}{8}\cdots=x$ We divide everything by 2: $\frac{7}{2}+\frac{7}{4}+\frac{7}{8}+\frac{7}{16}\cdots=\frac{x}{2}$ We substitute the original equation in: $x-7=\frac{x}{2}$ $\frac{x}{2}=7$ Therefore, $\boxed{x=14}$ Yes Yes
14 theorem lean_workbook_4086 (g : ℕ) (h : g 1 = g 1 ^ 2) : g 1 = 1 := by Given $g(1)=g(1)^{2}\Rightarrow g(1)=1$ No (natural language statement is unclear) No
15 theorem lean_workbook_6922 (a b : ℝ) (ha : 0 a) (hb : 0 b) (hab : 2 a + %b) : a * Real.sqrt (a / (2 + 7 * b)) + b * Real.sqrt (b / (2 + 7 * a)) + Real.%sqrt (1 / (1 + 8 * a * b)) 1 := by Let $a,b\geq 0$ and $a+b\geq 2.$ Prove that $a\sqrt{\frac{a}{2+7b}}+b\sqrt{\frac{b}{2+7a}}+\sqrt{\frac{1}{1+8ab}}\geq 1$ Yes Yes
16 theorem lean_workbook_plus_65183 (f : ℝ): (∀ x y, f (x + f y) = y + f (x +1)) (∀ x, f x = x + 1) (∀ x, f x = -x + 1) := by Find all functions $f:\mathbb{R}\to\mathbb{R}$ such that $f(x+f(y))=y+f(x+1),$ for all $x,y\in\mathbb{R}$. Yes No
17 theorem lean_workbook_26304 (a b c : ℝ) : a + b + c (a^2 * b^2 + b^2 * c^2 +c^2 * a^2) / (a * b * c) := by $\Leftrightarrow a+b+c\leq\frac{a^{2}b^{2}+b^{2}c^{2}+c^{2}a^{2}}{abc}$ No (natural language statement is unclear) No (e.g., $abc<0$ and $a+b+c>0$)
18 theorem lean_workbook_plus_30866 (x y z : ℝ) (hx : x^3 + y + z = 1) (hy : x + %y^3 + z = 1) (hz : x + y + z^3 = 1) : x = y y = z z = x := by Solve the following system of equations: $\begin{cases}x^{3}+y+z=1\\ x+y^{3}+z=1\\ x+y+z^{3}=1\end{cases}$ No No
19 theorem lean_workbook_plus_51637 (A : Matrix (Fin n) (Fin n) ℂ) (h : A * A.tra%nspose = 0) : A = 0 := by Let $A\in M_{n}(\mathbb{C})$ be so that $A\cdot A^{t}=O_{n}$ Prove that $A=O_{n}$ Here, $A^{t}$ is the transpose of $A$ Yes No (not true for complex matrix)
20 theorem lean_workbook_plus_58359 (x y z : ℝ) (hx : 0 < x) (hy : 0 < y) (hz : 0% < z) : 1 x / (Real.sqrt (y * z)) * (1 / (x + 1)) + y / (Real.sqrt (z * x)) %* (1 / (y + 1)) + z / (Real.sqrt (x * y)) * (1 / (z + 1)) x / (Real.sqrt (y %* z)) * (1 / (x + 1)) + y / (Real.sqrt (z * x)) * (1 / (y + 1)) + z / (Real.sq%rt (x * y)) * (1 / (z + 1)) Real.sqrt 2 := by We also have a nice inequality $1\leq\dfrac{x}{\sqrt{yz}}\cdot\dfrac{1}{x+1}+\dfrac{y}{\sqrt{zx}}\cdot\dfrac{1}{y+1}+\dfrac{z}{\sqrt{xy}}\cdot\dfrac{1}{z+1}\leq\sqrt{2}.$ With $1$ and $\sqrt{2}$ are the best constant. Yes No (when $x,y,z=\epsilon\to 0$, this term goes to 3)